OpenStudy (anonymous):

Write in standard form, an equation of the line passing through the given point with the given slope. slope = –8; (–2, –2)

OpenStudy (anonymous):

8x + y = –18 –8x + y = –18 8x – y = –18 8x + y = 18

OpenStudy (australopithecus):

use y=mx + b where, m = slope b = y-intercept (the y-intercept is the point where x = 0)

OpenStudy (australopithecus):

Just sub the slope that is given, and the x and y value of the point and solve for b. Remember a point in this case is defined as (x,y)

OpenStudy (australopithecus):

once you have solved for b just sub along with the slope into the equation y = mx + b remember the y value is the output and the x value is the input, so they are changing the slope and y-intercept are constants

OpenStudy (australopithecus):

any questions?

OpenStudy (anonymous):

So it would be B?

OpenStudy (australopithecus):

lol yes, wow they might as well have given you the answer

OpenStudy (anonymous):

Yay!

OpenStudy (australopithecus):

I hope you still go through the problem even though the answer is so evident to understand why it is correct

OpenStudy (anonymous):

I do and will. :)

OpenStudy (australopithecus):

-8(-2) + b = (-2) b = -16 - 2 b = -18 y = (-8)x + (-18)

OpenStudy (australopithecus):

y + 8x = -18

OpenStudy (australopithecus):

so no it is not b sorry I read the answers wrong ha I'm half asleep

OpenStudy (anonymous):

It happens to the best of us. :)

OpenStudy (australopithecus):

but yeah I hope you follow what I'm saying this stuff is important

OpenStudy (australopithecus):

even though it may not at the moment seem interesting or relevant